OpenStudy (anonymous):

Ms. Blackmon invested $8000, part at 4% and the rest at 5% per year. How much did she invest at each rate if her total return per year on the investments was $380? show work

OpenStudy (anonymous):

(8000 - x)0.04 + 0.05x = 380

OpenStudy (anonymous):

can u solve this equation

OpenStudy (anonymous):

my answer doesn't seem right

OpenStudy (anonymous):

x is one of the amounts he invested and as the total is 8000 the rest is invested at the other rate

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

x = 6000

OpenStudy (anonymous):

lets see what the eqwuation gives 8000 * 0.04 - 0.04x + 0.05x = 380 .01x = 380 - 320 = 60 x = 60 /.01 = 6000 thats right 6000 at 4percent and 2000 at 5 percent

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

thanks for your help!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I have another one posted if you can help